Lucas Company is considering investing in a new machine. The machine costs $14,300 and has an economic life of four years. The machine will generate cash flows of $4,200 (cash revenues less cash expenses) each year. All cash flows, except for the initial investment, are realized at the end of the year. The investment in the machine will be made at the beginning of the first year. Lucas is not subject to any taxes and, for financial accounting purposes, will depreciate the machine using straight-line depreciation over four years. Lucas uses a 10 percent cost of capital when evaluating investments. Use Exhibit A.9.

Required:

a. Calculate the accounting income for the total over four years.

b. Compute the NPV of the cash flows over four years. (Round PV factor to 3 decimal places. Negative amount should be indicated by a minus sign.)
Answer #1

a) Total Accounting income for 4 years = (4200-14300/4)*4 = $2500

b) Net present value = Present value of cash inflow-Present value of cash outflow

= (4200*3.169)-14300

Net present value = -990.20 or -990